OpenStudy (anonymous):

What is the midpoint between (-8, -5) and (-2, 2)?

OpenStudy (anonymous):

midpoint formula ( (x1+x2)/2, (y1+y2)/2) so, x = (-8-2)/2 = -5 y = (-5+2)/2 = -3/2 mid pt ( -5, -3/2)

OpenStudy (amistre64):

well, it would have been if i used the right numbers ... lol

OpenStudy (amistre64):